OpenStudy (anonymous):

5. The total cost of an item including sales tax is directly proportional to its price. If the total cost of a $24 item is $25.44, what is the total cost on a $75 item? A. $795 B. $79.50 C. $106 D. $76.44

OpenStudy (anonymous):

wont it be d?

OpenStudy (anonymous):

25.44/24 =x/75 75*25.44=24x x=75*25.44/24 79.5 or B